### v1

# Modules as effective nodes in coarse-grained networks of Kuramoto oscillators

Abstract
Most real-world networks exhibit a significant degree of modularity. Understanding the effects of such topology on dynamical processes is pivotal for advances in social and natural sciences. In this work we consider the dynamics of Kuramoto oscillators on modular networks and propose a simple coarse-graining procedure where modules are replaced by effective single oscillators. The method is inspired by EEG measurements, where very large groups of neurons under each electrode are interpreted as single nodes in a correlation network. We expose the interplay between intra-module and inter-module coupling strengths in keeping the coarse-graining process meaningful and show that its accuracy depends on the degree of intra-module synchronization. We show that, when modules are well synchronized, the phase transition from asynchronous to synchronous motion in networks with 2 and 3 modules is very well described by their respective reduced systems, regardless of the network structure connecting the modules. Application of the method to real networks with small modularity coefficients, on the other hand, reveals that the approximation is not accurate, although it still allows for the computation of the critical coupling and the qualitative behavior of the order parameter if the inter-module coupling is large enough.
I Introduction
Real world systems that exhibit synchronization are commonly contained in large networks of non-linear oscillators, such as neuronal networks [1, 2, 3, 4] and power grids [5, 6, 7, 8]. In these cases, measuring the individual state of each node is challenging, and corse-grained procedures are often employed [9]. For example, measurements of brain activity using functional magnetic resonance imaging (fMRI) [10, 11], near-infrared spectroscopy (NIRS) [12] and electroencephalograms (EEG) [13], assess the average behavior of large regions of the brain, instead of capturing the oscillations of individual neurons. Analysis of correlations between such regions is used to infer patterns related to resting states or the performance of specific tasks. Neuronal diseases such as autism, Parkinson, schizophrenia, and epilepsy, for example, are commonly associated with abnormal modular synchronization [14, 15, 16, 17, 18, 13].
Motivated by the difficulty of measuring and simulating large networks of oscillators, and by current EEG techniques, that always measures the average behavior of massive number of neurons in each electrode, we develop an approximation where groups of nodes are replaced by a single effective node, drastically reducing the size of the system. The theory is developed for modular networks, since modules form natural sub-groups of nodes, and we use the Kuramoto model as underlying dynamics.
In the Kuramoto model each oscillator is characterized by a single phase $\theta_{i}$ and its dynamics depends on other oscillators according to the equation [20, 19, 21, 22]
$\dot{\theta_{i}}=\omega_{i}+\frac{\lambda}{N}\sum_{j=1}^{N}\sin{(\theta_{j}-\theta_{i})}$ (1)
where $N$ is the total number of oscillators and the natural angular velocity $\omega_{i}$ is usually chosen from a symmetric and unimodal distribution $g(\omega)$ centered at $\omega_{0}$. Each oscillator interacts with all the others according to their phase difference and the interactions are modulated by a global coupling strength $\lambda$. In the limit $N\rightarrow\infty$, the system undergoes a continuous phase transition from disordered motion to synchronization at $\lambda_{\text{c}}=2/(\pi g(\omega_{0}))$. The phase transition can be characterized by the complex number
$z=re^{i\psi}=\frac{1}{N}\sum_{j=1}^{N}e^{i\theta_{j}}$ (2)
representing the average of all phases. The module of $z$ is the order parameter of the transition, going from $r=0$, when motion is disordered, to $r=1$ for perfect synchronization.
The original assumption that each oscillator interacts with all the others is a simplifying approximation that fails in many real systems, such as neurons in the brain, that are grouped in well-defined regions [23, 24, 25, 26, 27], and fireflies [28, 29, 30], that only interact with close neighbors. The extension of the Kuramoto model to networks [31, 32, 33, 34, 35] describing the set of possible pairwise interactions, revealed that the topology of connections has a large effect on the synchronization properties of the system, leading, for example, to frequency [36] and explosive synchronization [37, 38, 39, 40].
Here we propose an approximation to the dynamics of modular networks where each module is replaced by a single effective node. We compute the Kuramoto order parameter and study the conditions for the phase transition from disordered to synchronized state to be well represented by the reduced system. We find that such an approximation can be very good for a range of internal and inter-module coupling strengths. Moreover, we find that, under these conditions, the results are largely independent of the network of connections between modules, that can be ignored and replaced by a single link between the nodes representing the modules. This approach can be viewed as an alternative to other techniques based on the renormalization group that may rely on hyperbolic spaces [41], Laplacian matrices [42, 43, 44] and machine learning [45], which have been used to describe large networks in terms of fewer supernodes that retain the basic properties of the original system. Although our method only works well for networks with well defined modules, application to real networks lead to interesting, although not always accurate, approximations to the full dynamics. As examples we show numerical simulations for Zachary’s Karate club social network (two modules) and the C. Elegans gap junctions neural network (divided into three, five and ten modules).
This paper is organized as follows: in section II we develop our coarse graining procedure for modular networks and in sections III and IV we apply it to the simple cases of two and three modules respectively, using synthetic networks. In section V we apply the coarse graining method to Zachary’s Karate club and C. Elegans gap junctions networks. Finally in section VI we discuss our findings.
II Coarse graining for modular networks
Networks can be described by an adjacency matrix $A$ containing information about the coupling between pairs of nodes. Here we consider only undirected and binary matrices, where $A_{ij}=A_{ji}=1$ if nodes $i$ and $j$ interact and $A_{ij}=0$ otherwise. The extension of the Kuramoto model to networks is given by
$\dot{\theta_{i}}=\omega_{i}+\sum_{j=1}^{N}\lambda_{ij}A_{ij}\sin{(\theta_{j}-\theta_{i})}$ (3)
where $\lambda_{ij}$ defines the coupling strength between nodes $i$ and $j$.
It is usual in network theory to define modules as groups of nodes that are more densely connected to each other than to the rest of the network. Modular structures can be constructed by either decreasing the number of connections between nodes of different modules or by distinguishing between inner-module and outer-module connection strengths.
Here we construct synthetic modular networks as follows: modules are indexed by $\sigma=1,2,\dots,s$ and contain $N_{\sigma}$ nodes, or oscillators, with $N=\sum_{\sigma=1}^{s}N_{\sigma}$. We call $A^{\sigma\sigma^{\prime}}$ the block of the adjacency matrix between modules $\sigma$ and $\sigma^{\prime}$. We take pairs of oscillators that belong to the same module to be connected, i.e, $A_{ij}^{\sigma\sigma}=1$ for all $i,j\in\sigma$, and to have connection strength $\lambda_{\sigma\sigma}$. Nodes belonging to different modules, on the other hand, are connected with probability $p$ and strength $\lambda_{\sigma\sigma^{\prime}}=\lambda_{\sigma^{\prime}\sigma}$. Therefore, for $p=0$ the network has $s$ disjoint and fully connected modules, whereas for $p=1$ the modular character is manifested only by the different connection strengths.
For this type of network Eq.(3) can be written as
$\dot{\theta}_{\sigma,i}=\omega_{\sigma,i}+\sum_{\sigma^{\prime}=1}^{s}\frac{\lambda_{\sigma\sigma^{\prime}}}{\langle k\rangle_{\sigma\sigma^{\prime}}}\sum_{j=1}^{N_{\sigma^{\prime}}}A_{ij}^{\sigma\sigma^{\prime}}\sin{(\theta_{\sigma^{\prime},j}-\theta_{\sigma,i})}$ (4)
where $i=1,\dots,N_{\sigma}$ refers to an oscillator in module $\sigma$. Note that for a single fully connected module, the Kuramoto model, Eq. (1) has a normalization $1/N$ on the interaction term. Now that each oscillator has a varied number of connections, we changed $N$ to $\langle k\rangle_{\sigma\sigma^{\prime}}=\sum_{i}k_{i\sigma\sigma^{\prime}}/N_{\sigma}$, the average number of connections that an oscillator from module $\sigma$ has with the whole module $\sigma^{\prime}$, with $k_{i\sigma\sigma^{\prime}}=\sum_{j}A^{\sigma\sigma^{\prime}}_{ij}$. In the limit of a fully connected network we recover $\langle k\rangle_{\sigma\sigma^{\prime}}\rightarrow N_{\sigma^{\prime}}$. Note also that $\langle k\rangle_{\sigma\sigma^{\prime}}N_{\sigma}$ is the total number of connections between modules.
Summing Eq. (4) over $i$ and dividing by $N_{\sigma}$ we get
$\langle\dot{\theta}_{\sigma}\rangle=\langle\omega_{\sigma}\rangle+\sum_{\sigma^{\prime}=1}^{s}\frac{\lambda_{\sigma\sigma^{\prime}}}{\langle k\rangle_{\sigma\sigma^{\prime}}N_{\sigma}}\sum_{i=1}^{N_{\sigma}}\sum_{j=1}^{N_{\sigma^{\prime}}}A_{ij}^{\sigma\sigma^{\prime}}\sin{(\theta_{\sigma^{\prime},j}-\theta_{\sigma,i})}$ (5)
where $\langle\dot{\theta}_{\sigma}\rangle$ and $\langle\omega_{\sigma}\rangle$ denote the average velocity and natural frequency of module $\sigma$.
The average velocity is a variable regarding the module. If we were able to write Eq. (5) solely in terms of average module variables we could simplify the dynamics to $s$ equations, drastically reducing the dimensionality of the problem. This can be achieved in the special case where each and every module has a large enough internal synchrony such that the phases of oscillators belonging to the same module are about equal. In other words, if $\lambda_{\sigma\sigma}$ is large enough for $\theta_{\sigma,i}\approx\langle\theta_{\sigma}\rangle$, then we can lose the symbol $\langle\,.\,\rangle$ and arrive at
$\dot{\theta}_{\sigma}=\omega_{\sigma}+\sum_{\sigma^{\prime}=1}^{s}\lambda_{\sigma\sigma^{\prime}}\sin{(\theta_{\sigma^{\prime}}-\theta_{\sigma})}$ (6)
which is the Kuramoto model in a fully connected network, independent of the connection matrix $A_{ij}^{\sigma\sigma^{\prime}}$. Since the coefficients $\lambda_{\sigma\sigma^{\prime}}$ do not need to be all equal, this constitutes an asymmetric network: the strength of the interaction that module $\sigma$ has with module $\sigma^{\prime}$ is not necessarily the same as $\sigma^{\prime}$ has with $\sigma$.
Eq. (6) shows that, under model (4), the dynamics of $s$ modules is qualitatively the same as that of $s$ oscillators if high synchrony within modules is satisfied. This simple observation summarizes the analytical basis for the coarse-graining process. In the next sections we will quantify the validity of this result by simulating model (4) for $s=2$ and 3 to compare with the dynamics of 2 and 3 oscillators, respectively. Only symmetric cases will be treated, i.e, $\lambda_{\sigma\sigma}\equiv\lambda_{\text{in}}$ $\forall\sigma$ and $\lambda_{\sigma\sigma^{\prime}}=\lambda_{\sigma^{\prime}\sigma}\equiv\lambda/s$ $\forall\sigma\neq\sigma^{\prime}$. We shall see that Eq. (6) alone does not reproduce the behavior of the order parameter of the full modular network. However, when weights related to the degree of internal synchrony are introduced to renormalize the contribution of the coarse grained nodes, very good agreement is observed.
III Two modules
We start by briefly describing the dynamics of $N=2$ Kuramoto oscillators. In this case, Eq. (1) can be rewritten in terms of the oscillator’s phase difference $\phi\equiv\theta_{1}-\theta_{2}$ as
$\dot{\phi}=\omega-\lambda\sin{\phi}$ (7)
where $\omega\equiv\omega_{1}-\omega_{2}$ is the oscillator’s frequency offset. For $\lambda>\omega$, Eq. (7) has a stable fixed point at $\phi^{\ast}=\arcsin{(\omega/\lambda)}$ and the order parameter at the stationary solution can be written as
$r=\frac{1}{\sqrt{2}}\sqrt{1+\cos{\phi^{*}}}=\frac{1}{\sqrt{2}}\sqrt{1+\sqrt{1-\left(\frac{\omega}{\lambda}\right)^{2}}}.$ (8)
Note that when the system is not synchronized the order parameter can still be written as $r^{2}=(1+\cos{\phi})/2$, but now with a non stable phase offset. This allows us to write a mean and a deviation of $r$ for asynchronized regions. The results are $\text{E}[r]=2/\pi$ and $\text{Var}[r]=1/2-4/\pi^{2}$.
Now we consider a network with 2 modules. Since typical modules do not have the same size we will show simulations for the case where $N_{1}=200$ and $N_{2}=100$. The frequency distributions $g(\omega)$ are Gaussians of width $\Delta=1$ and centered at $\omega_{1}$ and $\omega_{2}=0$ respectively. In order to apply the coarse graining procedure we must first tune the coupling constants to guarantee that the oscillators in each module are sufficiently synchronized. We define the local order parameters
$z_{\sigma}=r_{\sigma}e^{i\psi_{\sigma}}=\frac{1}{N_{\sigma}}\sum_{j=1}^{N_{\sigma}}e^{i\theta_{\sigma,j}}$ (9)
and set minimum synchronization thresholds $q_{\sigma}$ for each module’s local order parameter. In the simulations we initially fixed $q_{\sigma}=0.9$ for both modules. This means that, for a given value of the inter-module coupling parameter $\lambda$ we need to find $\lambda_{in}$ such that $r_{\sigma}\geq q_{\sigma}$ for $\sigma=1,2$.
Figure 1 (a) illustrates the procedure for $\lambda=2$ on a network where the inter-module connection probability was set to $p=0.5$. We started by simulating Eq. (4) with internal coupling $\lambda_{\text{in}}=1$ for a fixed time interval of $\Delta t=8$. This assures the system went through the transient and stabilized. If, after this interval, $r_{\sigma}<q_{\sigma}$, we increase $\lambda_{\text{in}}$ by $0.1$ and continue the simulation for another $\Delta t$. We repeat the process until $r_{1}$ (green) and $r_{2}$ (yellow) surpass the boundary of $q_{\sigma}$ synchrony during a whole $\Delta t$. This determines the minimum $\lambda_{\text{in}}$ for the validity of the coarse graining procedure for given $\lambda$ and $p$.
Figure 1: (a): Numerical integration of Eq. (4) for $\lambda=2$. Green, yellow and purple curves are the order parameters $r_{1}$, $r_{2}$ and $r$ and dashed line represents $q_{\sigma}=0.9$ synchrony threshold (all referring to left $y$ axis). Black ladder function shows $\lambda_{\text{in}}$ values (referring to the right $y$ axis). (b)-(c): Final $\lambda_{\text{in}}$ and its respective $\langle r\rangle$ with one standard deviation $\sigma$ above and bellow in purple. In (b) magenta and turquoise lines marks the average of critical $\lambda$ and $\lambda_{\text{in}}$ over 3 simulations. In (c) triangles represent $\sigma>0.01$ and circles otherwise, green ribbon marks the 2 oscillators limits where the top green curve is Eq. (8) and bottom boundary is Eq. (12). Theoretical mean and standard deviation for the asynchronized region are shown with dashed green lines. Panel (d) shows the analogous of panel (c) for minimal modular synchrony of $q_{\sigma}=0.7$, lower ribbon boundary is given by Eq. (11) with $q_{\sigma}=0.7$. Simulation values: $N_{1}=200$, $N_{2}=100$ and $p=0.5$. Gaussian frequency distributions with $\Delta=1$ and $(\omega_{1},\omega_{2})=(1.5,0)$.
To compare the dynamics of the modular system to that of two oscillators, however, requires information on global synchronization. The purple curves in Figures 1 (a) represents the global order parameter. After reaching minimum $\lambda_{\text{in}}$ for internal synchrony we let the system evolve for a new time interval of $\Delta t=20$. This allows us to calculate the average and standard deviation of $r$ and, consequently, classify the state of the full system as synchronized or not. Note the stable value of $r$ indicating the small standard deviation.
Calculating the standard deviation for a set of values of $\lambda$, instead of just the one presented in Figure 1 (a), we get panel (c) showing the average order parameter $\langle r\rangle$ as well as one unit of standard deviation $\sigma$ after reaching minimum $\lambda_{\text{in}}$, both in purple. For small $\lambda$, $r$ fluctuates and $\sigma$ is large, whereas for large $\lambda$ the order parameter converges to a stationary value with small standard deviation. We define the phase transition to global synchrony as the lowest $\lambda$ such that $\sigma$ is lower than the threshold of $1\%$. In the figure we use triangles for non-synchronized states and dots for synchronized states.
Fitting the global synchronization of two modules by Eq. (8) puts and upper boundary for the fit, shown as the upper green curve of Figure 1 (c). This happens because we force modules to have $q_{\sigma}\leq r_{\sigma}\leq 1$. Given that we slowly increase $\lambda_{\text{in}}$, a lower boundary for $r(\lambda)$ would constitute a better fit for the synchrony. To derive such equation we modify Eq. (2) and calculate the global order parameter taking into account the weights $q_{\sigma}$, related to the internal synchrony of each module:
$z=re^{i\psi}=\frac{1}{s}\sum_{\sigma=1}^{s}q_{\sigma}e^{i\theta_{\sigma}}.$ (10)
The weights $q_{\sigma}$ compensate for the fact that the module is not in perfect synchrony. For two modules with weights $q_{1}$ and $q_{2}$ we find
$r=\frac{1}{2}\sqrt{q_{1}^{2}+q_{2}^{2}+2q_{1}q_{2}\cos{\phi}}.$ (11)
which should replace Eq. (8).
Plugging in the minimum synchrony $q_{\sigma}=0.9$ for both modules we obtain
$r=\frac{0.9}{\sqrt{2}}\sqrt{1+\sqrt{1-\left(\frac{\omega}{\lambda}\right)^{2}}}.$ (12)
This is the lower bound curve shown in green in Figure 1 (c). The green ribbon marks the region between the upper and lower boundaries.
Note that it is impossible for the modules to synchronize at values above the upper boundary because it is a barrier of perfect oscillators which our modules would only asymptomatically reach when $\lambda_{\text{in}}\rightarrow\infty$. This also means that the synchronization can fall bellow the lower boundary as we see for some points of Figure 1 (c). Nonetheless, Eq. (12) constitutes a good fit for the global synchrony of the system. Going the opposite way and loosening our condition on minimum module synchrony to $r_{\sigma}\geq q_{\sigma}=0.7$ we get Figure 1 (d). The simulated values fall, mostly, within the ribbon which is clearly bigger due to the lower boundary now being Eq. (11) with $q_{1}=q_{2}=0.7$. Even though we can theoretically define the green ribbon as a reasonable region for the predicted global synchronization, the looser we let modules be the bigger is the region and consequently less insightful it is.
Recall that for each $\lambda$ we had to determine a minimum $\lambda_{\text{in}}$ that fulfills the condition $r_{\sigma}\geq 0.9$. These values are shown in Figure 1 (b). We obtain a tent shaped curve that creates two regions: above the curve the coarse graining process is valid, and bellow the curve, it is not. Global phase transition is marked by the magenta line, while turquoise indicates the respective $\lambda_{\text{in}}$ for this globally synchronized state. The onset of synchronization is marked by the intersection of the magenta and turquoise lines.
Observe that starting from global asynchrony, it gets harder to satisfy $r_{\sigma}\geq 0.9$ as we increase $\lambda$ thus demanding larger $\lambda_{\text{in}}$, i.e, the strength of connections among modules disrupts the cohesion within the modules. On the other hand, after transitioning to global synchrony the parameter $\lambda$ helps bringing modules close together and consequently, demands less from the inner-module coupling strength.
The onset of synchronization indicates both minimum $\lambda_{\text{in}}$ for $r_{\sigma}\geq 0.9$ and $\lambda$ for global synchrony. Figure 2 shows such $(\lambda_{\text{in}},\lambda)$ pairs for 5 different values of frequency offsets $\omega$ and 5 different inter-module connection probabilities $p$. Panel (a) shows that, for large enough $\lambda_{\text{in}}$, the global phase transition happens at about $\lambda=\omega$ which is precisely when the phase transition of a 2 oscillator system takes place, in accordance with the coarse-graining process. More interesting, the transition point does not depend upon the number of connections between modules, as expected from Eq. (6). Panel (b) shows that immediately after the transition to the globally synchronized state, the minimum $\lambda_{\text{in}}$ that would assure sufficient synchrony for the coarse-graining process decreases non linearly with the amount of connections between the modules, exposing the importance of network structure in keeping inner-module cohesion.
Figure 2: Dots represents the numerical integration of system (4) for 3 different connection probabilities $p$ and frequency offsets $\omega$. (a): Critical inter-module coupling strength. Lines represent a 2 oscillators system. (b): Minimum inner-module coupling strength at phase transition. Lines simply connect the points. Global values: $N_{1}=200$ and $N_{2}=100$. Gaussian frequency distributions with $\Delta=1$ and $(\omega_{1},\omega_{2})=(1.5,0)$. Error bars made over 3 simulations.
IV Three modules
We now extend the results of the previous section to systems with three modules. We first derive an expression for $r$ when $N=3$. The Kuramoto dynamics can be written in terms of the phase differences $\phi_{12}\equiv\theta_{1}-\theta_{2}$ and $\phi_{23}\equiv\theta_{2}-\theta_{3}$ as
$\dot{\phi}_{12}$ $=\Omega_{1}-2\sin{(\phi_{12})}-\sin{(\phi_{12}+\phi_{23})}+\sin{(\phi_{23})}$ (13a)
$\dot{\phi}_{23}$ $=-\Omega_{3}+\sin{(\phi_{12})}-\sin{(\phi_{12}+\phi_{23})}-2\sin{(\phi_{23})}$ (13b)
where we rescaled time $t\rightarrow 3t/\lambda$, set $\omega_{2}=0$ (which is equivalent to a change in reference frame) and defined $\Omega_{1}\equiv 3\omega_{1}/\lambda$ and $\Omega_{3}\equiv 3\omega_{3}/\lambda$. These equations are similar to Adler’s equation $\dot{\phi}=\Omega-\sin\phi$, but in 2 dimensions. Note that every point in the $\phi_{12}\times\phi_{23}$ plane is a fixed point for some set of parameters $\{\Omega_{1},\Omega_{3}\}$, nonetheless, only some of these fixed points will be stable. Fixed points are solutions of
$\Omega_{1}=2\sin{(\phi_{12})}+\sin{(\phi_{12}+\phi_{23})}-\sin{(\phi_{23})}$ (14a)
$\Omega_{3}=\sin{(\phi_{12})}-\sin{(\phi_{12}+\phi_{23})}-2\sin{(\phi_{23})}.$ (14b)
Similar to Eq. (8), the order parameter can be written in terms of phase differences as
$r=\frac{1}{3}\sqrt{3+2[\cos{\phi_{12}}+\cos{(\phi_{12}+\phi_{23})}+\cos{\phi_{23}}]}.$ (15)
Figure 3 (a) shows the contour plot of Eq. (15) in the phase space $\phi_{12}\times\phi_{23}$. To determine the stability of these fixed points we calculated the Jacobian of system (13) and looked at the largest real part of the pair of eigenvalues: when it is zero we are at the critical line, shown as the red curve in Fig.3 (a) . The curve divides the phase space into two regions, inside where the system displays stable synchrony, and outside which presents periodic motion. More elucidating, however, is to understand synchronization in the parameter space $\Omega_{1}\times\Omega_{3}$. Figure 3 (b) shows the stable $r$ values from panel (a) reparameterized according to Eq. (14). It is very interesting to note that, regardless of being in the phase space or the parameter space, contours of constant synchronization resemble ellipses. We proceed to finding approximations for such contours. This will help fit simulations from three-module networks with the three-oscillator system.
Figure 3: (a): Eq. (15) in the phase space $\phi_{12}\times\phi_{23}$. Red curve divides stable from unstable points, stable within. (b): Stable points reparameterized to the parameter space $\Omega_{1}\times\Omega_{3}$ via Eq. (14). (c): How stable $r$ values changes through the ellipses. Dashed lines are the empirical values from Eq. (18). (d): Contrast between the ellipses (17) (dashed) and constant $r$ values (continuous).
In a simpler version of the problem, if any pair of the 3 oscillators is identical, stable solutions will imply equal phases for such pair. This means that either $\phi_{12}$, $\phi_{23}$ or $\phi_{12}+\phi_{23}$ is zero depending on the pair that is identical. If the remaining phase difference also reaches a fixed point, we can determine it through system (13) and replace it into Eq. (15) to write $r$ in terms of $\Omega$, the frequency difference between the non-identical oscillators, as
$r=\frac{1}{3}\sqrt{5+4\sqrt{1-\frac{\Omega^{2}}{9}}}.$ (16)
Eq. (16) gives the synchronization order parameter when 2 oscillators are identical, i.e, along the axes $\Omega_{1}=0$ $(\Omega=\Omega_{3})$, $\Omega_{3}=0$ $(\Omega=\Omega_{1})$ and $\Omega_{1}=\Omega_{3}$ $(\Omega=\Omega_{1}=\Omega_{3})$ of Figure 3 (b). Thus, we empirically propose the constant synchronization contours to be ellipses described by
$\Omega_{1}^{2}-\Omega_{1}\Omega_{3}+\Omega_{3}^{2}=\alpha$ (17)
and we modify Eq. (16) into
$r=\frac{1}{3}\sqrt{5+4\sqrt{1-\frac{\alpha}{9}}}$ (18)
with $\alpha\leq 9$ for the general case. This approximation indicates a first order phase transition from periodic motion to synchronization at $\alpha\equiv\alpha_{1}=9$ or, in terms of the original parameters $\omega_{1}$, $\omega_{3}$ and $\lambda$, at $\lambda_{\text{c}}^{2}=\omega_{1}^{2}-\omega_{1}\omega_{3}+\omega_{3}^{2}$.
Figures 3 (c)-(d) show the accuracy of this expression. The closer $(\Omega_{1},\Omega_{3})$ is to the origin, and consequently the lower the $\alpha$, the better the ellipse matches the real constant $r$ contour. As we move away from the origin, our proposed ellipses start to cut through several real fixed $r$ contour lines. In that case, what we observe is that Eq. (18) is returning the lowest synchronization value attainable over all the real contours that it is crossing.
In the inset of Figure 3 (d) we can easily see that there are constant $r$ values that fall out of the theoretical ellipses, therefore, there should be a smaller ellipse that encompass all values. We numerically find that further increasing $\alpha$ up to $\alpha_{2}\equiv 9.295$ causes all possible synchronization outcomes to tightly fall within the ellipse. In other words, depending on how we control our parameters $\{\Omega_{1},\Omega_{2}\}$ the phase transition happens somewhere between $\alpha_{1}\leq\alpha\leq\alpha_{2}$, or in terms of the original parameters, $\lambda^{2}\leq\omega_{1}^{2}-\omega_{1}\omega_{3}+\omega_{3}^{2}\leq 1.0328\lambda^{2}$. Still, we have no approximation for $r$ in such regime.
Figure 4: (a)-(c): Final $\lambda_{\text{in}}$, global order parameter $r$ and standard deviation $\sigma$ in the frequency parameter space. Simulation values: $N_{1}=90$, $N_{2}=100$, $N_{3}=110$, $p=0.01$ and $\lambda=3$. Gaussian frequency distributions with $\Delta=1$ and means $\omega_{1}$, $\omega_{2}=0$ and $\omega_{3}$.
We now proceed to simulate Eq. (4) with three modules, as before we take modules of distinct sizes to be more realistic, namely $N_{1}=90$, $N_{2}=100$ and $N_{3}=110$. Throughout this section we fixed $p=0.01$ and $\lambda=3$ so that $\Omega_{i}=\omega_{i}$. Again, we set the frequency distributions to be Gaussians of width $\Delta=1$ and centers $\omega_{1}$, $\omega_{2}=0$ and $\omega_{3}$. The minium synchrony of each module is $q_{\sigma}=0.9$.
As before, we calculate the standard deviation of $r$. Figures 4 (b)-(c) respectively show the final value of the global order parameter and its standard deviation. It is also plotted in black the elliptical boundary of Eq. (17). We observe that the synchronization region corresponds to the expected behavior for three oscillators and exhibit all distinct behaviors, full and partial phase locking and asynchrony.
For three modules we have seen so far that given high enough $\lambda_{\text{in}}$ the coarse-graining procedure retain the basic properties of the reduced system; now we discuss how different parameter values affect the internal coupling. Figure 4 (a) shows the final value of $\lambda_{\text{in}}$. We see the same behavior of two modules, approaching the synchronization regime from a asynchronous state increases the internal coupling required to preserve $r_{\sigma}\geq q_{\sigma}$. This might seem an odd behavior considering we are decreasing the absolute frequency value and consequently expect it to be easier to keep cohesive modules. However, given that there are few connections between modules, the transfer of information from one to the others is harder when their frequencies are further apart, even though the absolute value is smaller.
V Real Networks
So far we have seen conditions for the known behavior of 2 and 3 oscillators emerge in artificial modular networks. These networks, however, were generated in such a way as to exhibit very strong modular structure that does not resemble real world systems. In what follows we test our results on two famous networks, Zachary’s Karate club social network and the C. Elegans gap junctions neural network [46, 26]. Both of these networks can be modularized. What distinguishes them from the artificial networks constructed in this paper are structural properties such as not having full number of connections within modules and non random connections between modules per se. We show the effects of the coarse-graining procedure in these networks.
Figure 5: Top: Dynamics of Eq. (4) for Zachary’s Karate social network. (a): Final $\lambda_{\text{in}}$. Magenta and turquoise lines marks the average of critical $\lambda$ and $\lambda_{\text{in}}$. (b): $\langle r\rangle$ with $\lambda_{\text{in}}$ relative to (a) in purple and with fixed $\lambda_{\text{in}}=20$ in red, green ribbon represents Eq. (11) for $0.9\leq q_{\sigma}\leq 1$. Simulation values: $\Delta=0$, $\omega_{1}=1.5$ and $\omega_{2}=0.0$. Bottom: Dynamics of Eq. (4) for C. Elegans gap junctions network. (c)-(e): $\langle r\rangle$ for the network with fixed $\lambda_{\text{in}}=20$ in red, equivalent coarsed-grained system in green. Simulation values: $\Delta=0$, (a): $(\omega_{1},\omega_{2},\omega_{3})=(0,1,6)$; (b): $(\omega_{1},\omega_{2},\omega_{3},\omega_{4},\omega_{5})=(0,1,1,-2,7)$; (c): $(\omega_{1},\omega_{2},\omega_{3},\omega_{4},\omega_{5},\omega_{6},\omega_{7},\omega_{8},\omega_{9},\omega_{10})=(0,1,-3.5,-2,2,0.6,0.4,1.2,1.5,-1.5)$.
Let us start with the Karate network [46]. It has a total of 34 nodes with 78 connections divided into two modules of 17 nodes each. As before, we put this network under the dynamics given by Eq. (4) and slowly increase $\lambda_{\text{in}}$ until a minimum inner module synchrony of 90% is reached. Figure 5 (a) shows the minimum internal coupling that keeps the internal synchrony threshold immediately after the phase transition. As expected, increasing $\lambda$ for asynchronized states impairs modular synchrony, hence the continuous increase in $\lambda_{\text{in}}$ as $\lambda$ grows. On the other hand, after the phase transition, synchronized states act in cohesion, requiring lesser $\lambda_{\text{in}}$ as we increase $\lambda$. Note that the phase transition is at the expected value of 1.5 and thus is not affected by the structural differences of Zachary’s network. However, when comparing Zachary’s network to the artificial network used in the 2 modules section, synchronized states do not reach the synchronization values of 2 oscillators system, this can be seen in Figure 5 (b) where the purple curve shows the $r$ values for Zachary’s network, note that they do not fall within the 2 oscillators ribbon shown in green. Therefore, the structural differences between the real and artificial networks are affecting the coarse-graining procedure. We can counteract the effects of network structure by overshooting the internal coupling, the red curve of Figure 5 (b) shows the synchronization values when keeping $\lambda_{\text{in}}=20$ throughout, this forces the system to behave as 2 oscillators in spite of networks effects.
As for our second example we study a reduced network from the C. Elegans gap junctions neural network containing 248 nodes and 511 connections. Using appropriate metrics it can be modularized into 3, 5 or 10 modules [47, 48]. Note that regardless of number of modules, some end up being far greater than others, for instance, in the 3 modules arrangement we have 130 nodes in the largest module, 77 nodes in the intermediate and 41 nodes in the smallest, for 5 modules the number of nodes in each module are 120, 83, 34, 7 and 4 and with 10 modules the range goes from 76 nodes to only 6. Add to that the fact that the network has about 1% of the total number of possible connections to conclude that it is very hard to make all modules have a minimum synchrony of 90% or even lower. For that reason we choose only to overshoot $\lambda_{\text{in}}$ and compare whether the global order parameter of the network resembles that of the coarse-grained system for the cases of 3, 5 and 10 oscillators. Figure 5 (c)-(e) shows in red the simulation of $\langle r\rangle$ for those 3 modularizations, each accompanied by a simulation of the order parameter of the reduced system in green for comparison. For 3 modules results show that despite the lack of oscillatory behavior by the neural network, exposed by the low standard deviation ribbon, synchronization values fall somewhat in the mean values of the coarse-grained system, the same is valid for 10 modules. Differently, 5 modules achieve global synchronization much faster in the neural network then in the reduced system, this is most likely due to the modularization process. As mentioned earlier, the C. Elegans gap junctions neural network is biased, not all modules are equally sized, however, specifically for 5 modules the discrepancy in amount of nodes in each group is much higher then in the other settings, this causes the system to be highly dependent on very specific modules which, if synchronized enough trough $\lambda_{\text{in}}$ values, already gives global synchronization regardless of other modules.
VI Conclusions
In this work we studied a coarse-graining process in modular networks where each module is reduced to a single effective node. We constructed artificial networks where modules are fully connected and nodes interact with strength $\lambda_{\text{in}}$ for intra-modular connections, and $\lambda$ for inter-module connections, that exist with probability $p$. Using Kuramoto oscillators as underlying dynamics we derived the Eq. (6) for the average dynamics of the modules, which is independent of the network details. This constitutes the base of our coarse-graining process, leading to the approximation where the dynamics of $s$ modules would be equivalent to that of a system composed of $s$ oscillators. We tested this prediction by simulating the Kuramoto model with 2 and 3 modules. Moreover, we included the assumption of high internal synchrony into the estimate of the order parameters via $r_{\sigma}\geq q_{\sigma}$. Thus, our coarse-graining procedure gives a lower bound for the global network behaviors when the system is replaced by $s$ oscillators, where $q_{\sigma}$ is the minimum modular synchrony.
Results of the coarse-graining method are shown in Figures 1 and 4 for 2 and 3 modules. Figures 1(b) and 4(a) show that, starting from global asynchrony but keeping $r_{\sigma}\geq q_{\sigma}$, approaching the first order phase transition demands increasingly larger inner-module coupling strength. In this regime the connections among the modules, responsible for bringing them together, hinder internal module’s synchrony. On the other hand, after the phase transition, the inter-module connections facilitate global and internal synchrony, as they bring modules close together. Figures 2(a) and 4(b)-(c) show that the parameter values at which the system transitions is the same as the equivalent $s$ oscillators system, validating the coarse-graining method. Nonetheless, assuming high modular synchrony as we did is a strong assumption because it is highly dependent upon the number of connections between modules, as we see from Figure 2 (b), where fewer connections creates a demand for higher $\lambda_{\text{in}}$. We naturally expect this structural role to also be contained in inner-modular connections, with more connections within modules positively correlating to lower internal coupling.
We also applied the method to real networks where modules are not so well defined. We considered a small network with two modules, Zachary’s Karate club social network, and the C. Elegans gap junctions neural network, which can be modularized in different ways. Our results, displayed in Figure 5, show that the complexity these networks introduces difficulties that cannot be captured by the simple coarse-graining processes proposed here. The reduction of whole modules into single oscillators is not sufficient to describe the synchronization process in detail. However, some intuition can still be gained by the method, such as the approximate value of critical coupling and even the behavior of the order parameter if the internal coupling is large enough. Yet, the inaccuracy of the approximation raises the question of how reliable are the dynamics inferred from networks obtained from measurements of brain activity, e.g., where large areas of the brain, not necessarily comprising a module, are replaced by a single node, such as in functional magnetic resonance imaging, near-infrared spectroscopy and electroencephalograms.
Network properties like size, topology and connectivity are important to explain the disparity in synchronization between real and coarse-grained networks. Zachary?s network e.g. has higher connectivity to number of nodes ratio than C. Elegans network. It has also been shown that to retain synchronization, some Laplacian eigenvalues should be preserved between the original and coarse-grained networks [49]. Understanding the role of these properties in the coarse-graining process will help to unravel under what circumstances treating groups of nodes as single oscillators is a valid assumption.
Acknowledgements.
It is a pleasure to thank J. A. Brum and Guilherme F. Arruda for helpful suggestions. This work was partially supported by FAPESP, grant 2021/14335-0 (MAMA) CAPES, grant 88887.706116/2022-00 (LLB), and CNPq, grant 303814/2023-3 (MAMA).